40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3936     Entry time: Tue Nov 16 23:36:29 2010
Author: Suresh, Jenne 
Type: Update 
Category: SUS 
Subject: Assembly of ETMs 

[Jenne, Suresh]

 

The ETM assembly has moved forward a couple of steps.  We have completed the following:

1) Positioning the guide rod and wire stand-off on both the ETMs (5 and 7)

2) The magnets had to be cleaned with an acetone wash as they had touched the plastic Petri-dish (not cleaned for vacuum).

3) The magnets and the Al dumb-bells have been glued together and left to cure in the gluing fixture.

4) The guide-rod and wire stand-offs have also been glued to the optic and left to cure for 24 hrs.

 

 

JD:  As you can see in my nifty status table, we are nearing the end of the suspension story.  

StatusTable.png

We are going to try (but can't guarantee) to get ETMX to Bob for baking by Friday at lunchtime, that way we can re-suspend it on ~Monday, and place it in the chamber.  Then we could potentially begin Green arm locking next week.  Steve has (hopefully!!) ordered the spring plungers for ETMY.  The receiving and baking of the spring plungers is the only current delay that I can foresee, and that only is relevant for one of the optics. 

We (who is going to be in charge of this?) still need to move the SRM OSEMs & cables & connectors to the ITMY chamber from the BS chamber. 

 

 

ELOG V3.1.3-